OpenStudy (anonymous):

Graph the line with slope -2 passing trough the point (3,5)

OpenStudy (anonymous):

To find the equation, you would use point slope form, because you have the slope and the point. Point slope form is Y-Y1=m(X-X1). M=-2 (m is your slope), X1=3, and Y1=5. Now you can write your equation. It would look like Y-5=-2(X-3). Distribute your -2 to entering inside the parentheses. Now your equation is Y-5=-2X+6. You want Y to be alone on its side, so add 5 to both sides. Adding 5 to the left side would cancel out your -5, and you would combine 5 with 6 on the right side. Your solved equation should look like Y=-2X+11. Now just graph and you're done :)

OpenStudy (anonymous):

Sorry, everything, not entering.